OpenStudy (anonymous):

Robin puts 5/8 cups of water into her recipe. Noel puts 1/3 the amount of water that Robin puts in her recipe. How much water does Noel put in her recipe?

OpenStudy (anonymous):

So what you want to do is multiply \[\frac{ 5 }{ 8 } * \frac{ 1 }{ 3 } \to \frac{5}{24}\] That means that Noel put in \[\frac{5}{24} \] cups of water in her recipe.

OpenStudy (anonymous):

ok thank you